# Trivial Isomorphisms between Reduced Products

### v5

@@ 5.2. Uniformization
Assume $\operatorname{{\mathsf{OCA}_{T}}}$. If ${\mathcal{M}}_{n}$ and ${\mathcal{N}}_{n}$, for $n\in{\mathbb{N}}$, are countable sets and
@@
is a coordinate respecting function with $\alpha^{\Phi}=\operatorname{id}$ such that the ideal ${\mathcal{J}}_{\rm{prod}}$ is nonmeager, then $\Phi$ has a Borel-measurable lifting.
@@
Following the notation set up before Definition 2.6, if $A\in\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$, then $\Phi_{A}$ is the induced function between $\prod_{n}{\mathcal{M}}_{n}/\operatorname{Fin}\restriction A$ and $\prod_{n\in A}{\mathcal{N}}_{n}/\operatorname{Fin}\restriction A$.
@@
of functions with $g^{A}_{n}\colon{\mathcal{M}}_{n}\to{\mathcal{N}}_{n}$ such that the function $\Psi_{A}:\prod_{n\in A}{\mathcal{M}}_{n}\to\prod_{n\in A}{\mathcal{N}}_{n}$ defined by
@@
Consider the Polish space ${\mathcal{Y}}_{n}=({\mathcal{N}}_{n})^{{\mathcal{M}}_{n}}$. Then $\{(g^{A},A)\mid A\in{\mathcal{J}}_{\rm{prod}}\}$ is a family of partial selectors as defined and studied in §3.1. The following shows that the functions $g^{A}$, for $A\in{\mathcal{J}}_{\rm{prod}}$, moreover form a coherent family in the sense of [9, §2].
@@
Proposition 5.2 implies that ${\mathcal{J}}_{\rm{prod}}$ is nonmeager (recall that this conclusion is included in the assumptions of Proposition 5.6), and (again by Talagrand and Jalali-Naini, [9, Theorem 3.10.1]) it intersects every perfect tree-like almost disjoint family. Fix $B\in{\mathcal{A}}\cap{\mathcal{J}}_{\rm{prod}}$. For $x\in Z$ let $n(x)=\max(\operatorname{Diff}(g^{f(x)},g^{B}))$ (with $\max\emptyset=0$). Since $Z$ is uncountable, there is $n$ such that $Z^{\prime}=\{x\in Z\mid n(x)<n\}$ is uncountable. Choose distinct $y,z$ in $Z^{\prime}$ satisfying $\Delta(y,z)\geq n$ and fix $j$ in $(\operatorname{Diff}(g^{f(y)},g^{f(z)})\cap B)\setminus n$. Then $g^{B}_{j}=g^{f(y)}_{j}\neq g^{f(z)}_{j}=g^{B}_{j}$; contradiction. ∎
Since ${\mathcal{J}}_{\rm{prod}}$ is nonmeager, there is $\bar{n}$ such that ${\mathcal{G}}_{\bar{n}}$ as provided by Claim 5.8 is nonmeager. Consider the hereditary closure of ${\mathcal{G}}_{\bar{n}}$,
$\hat{\mathcal{G}}_{\bar{n}}=\{A\subseteq{\mathbb{N}}\mid(\exists B\in{\mathcal{G}}_{\bar{n}})A\subseteq B\}.$
[9, Theorem 3.10.1(a)] implies that there exists $\bar{k}$ such that every finite subset of ${\mathbb{N}}\setminus\bar{k}$ is in $\hat{\mathcal{G}}_{\bar{n}}$. Let
${\mathcal{Y}}=\{(a,b)\in(\prod_{n}{\mathcal{M}}_{n})\times(\prod_{n}{\mathcal{N}}_{n})\mid(\forall m\geq\bar{k})(\exists X\in{\mathcal{G}}_{\bar{n}})\ [\bar{k},m)\subseteq X\\ \text{ and }b(i)=g^{X}_{i}(a(i))\text{ for all $\bar{k}\leq i<m$}\}.$
Clearly ${\mathcal{Y}}$ is a closed subset of $(\prod_{n}{\mathcal{M}}_{n})\times(\prod_{n}{\mathcal{N}}_{n})$.
@@
If $(a,b)\in{\mathcal{Y}}$, then $b=^{*}\Phi_{*}(a)$.
@@
Assume otherwise and fix $(a,b)\in{\mathcal{Y}}$ such that $S=\{n\mid b(n)\neq\Phi_{*}(a)(n)\}$ is infinite. Since $\hat{\mathcal{G}}_{\bar{n}}$ is nonmeager and hereditary, there is an infinite $X\subseteq S$ in $\hat{\mathcal{G}}_{\bar{n}}$ ([9, Theorem 3.10.1 (a)] gives a considerably stronger statement). Let $Y\in{\mathcal{G}}_{\bar{n}}$ be such that $X\subseteq Y$ and note that $\Psi_{Y}(a)=^{*}\Phi_{*}(a)\restriction Y$. Let $l\geq\bar{k}$ be such that $g^{Y}_{i}(a(i))=\Phi_{*}(a)(i)$ for all $i\geq l$ and find $m$ such that $X_{0}=X\cap[l,m)$ has at least $\bar{n}+1$ elements. By the definition of ${\mathcal{Y}}$, there is $C\in{\mathcal{G}}_{\bar{n}}$ such that $g^{C}_{i}(a(i))=b(i)$ for all $i\in X_{0}$. This implies $\Delta(g^{C},g^{Y})\supseteq X_{0}$. Since $|X_{0}|\geq\bar{n}+1$, this contradicts the homogeneity property of ${\mathcal{G}}_{\bar{n}}$ as in Claim 5.8. ∎
Claim 5.9 implies that for every $a\in\prod_{n}{\mathcal{M}}_{n}$ all elements of the vertical section
${\mathcal{Y}}_{a}=\{b\in\prod_{n}{\mathcal{N}}_{n}\mid(a,b)\in{\mathcal{Y}}\}$
are equal to $\Phi_{*}(a)$ modulo finite, in particular ${\mathcal{Y}}_{a}$ is countable. By the Luzin–Novikov Theorem ([27, Theorem 18.10]), there is a Borel-measurable uniformization $\Theta\colon\prod_{n}{\mathcal{M}}_{n}\to\prod_{n}{\mathcal{N}}_{n}$. Claim 5.9 implies that $\Theta$ is a lifting of $\Phi$. ∎
5.3. Automorphisms with Baire-measurable liftings
Proposition 5.10.
Assume $\operatorname{{\mathsf{OCA}_{T}}}$. If ${\mathcal{M}}_{n}$ and ${\mathcal{N}}_{n}$, for $n\in{\mathbb{N}}$, are countable sets and
is a coordinate respecting function with $\alpha^{\Phi}=\operatorname{id}$ such that $\Phi$ has a Baire-measurable lifting, then $\Phi$ is of product form.
Proof.
Let $\Theta\colon\prod_{n}{\mathcal{M}}_{n}\to\prod_{n}{\mathcal{N}}_{n}$ be a Baire-measurable function which is a lifting for $\Phi$. We want to construct functions $h_{n}\colon{\mathcal{M}}_{n}\to{\mathcal{N}}_{n}$ such that for all $a\in\prod_{n}{\mathcal{M}}_{n}$, for all but finitely many $n\in{\mathbb{N}}$, $\Theta(a)(n)=h_{n}(a(n))$.
Consider first the case when all ${\mathcal{M}}_{n}$ are finite. In this case we will not need $\operatorname{{\mathsf{OCA}_{T}}}$. Since $\Theta$ is Baire-measurable, it is continuous on a dense $G_{\delta}$ set $G$. By the standard representation of dense $G_{\delta}$ sets in $\prod_{n}{\mathcal{M}}_{n}$ (see e.g., [13, Theorem 9.9.1]) there are a partition ${\mathbb{N}}=S_{0}\sqcup S_{1}$, $a_{0}\in\prod_{n\in S_{0}}{\mathcal{M}}_{n}$, and $a_{1}\in\prod_{n\in S_{1}}{\mathcal{M}}_{n}$ such that for $j\in\{0,1\}$ we have $\{b\in\prod_{n}{\mathcal{M}}_{n}\mid b\restriction S_{j}=a_{j}\}\subseteq G$. (This is precisely where the finiteness of ${\mathcal{M}}_{n}$ gets used, as this characterisation does not hold for infinite sets). If $x\in\prod_{n\in S_{0}}{\mathcal{M}}_{n}$ and $y\in\prod_{n\in S_{1}}{\mathcal{M}}_{n}$, let $x^{\frown}y$ be the concatenation of $x$ and $y$, that is, the unique $z\in\prod_{n}{\mathcal{M}}_{n}$ which equals $x$ on $S_{0}$ and $y$ on $S_{1}$.
Therefore the function $\Upsilon\colon\prod_{n}{\mathcal{M}}_{n}\to\prod_{n}{\mathcal{N}}_{n}$ defined by
$\Upsilon(b)=(\Theta(b\restriction S_{0}^{\frown}a_{1})\restriction S_{0})^{\frown}(\Theta(b\restriction S_{1}^{\frown}a_{0})\restriction S_{1})$
is continuous and $\Upsilon(b)=^{*}\Theta(b)$ for all $b\in\prod_{n}{\mathcal{M}}_{n}$.
For $m\in{\mathbb{N}}$ and $u\in\prod_{j<m}{\mathcal{M}}_{j}$ write
$[u]=\{a\in\prod_{n}{\mathcal{M}}_{n}\mid a\text{ extends }u\}.$
Extend $\Upsilon$ to a partial function on $\bigcup_{m}\prod_{j<m}{\mathcal{M}}_{j}$ by setting (keep in mind that the continuity of $\Upsilon$ reduces to the fact that $a\mapsto\Upsilon(a)(n)$ is locally constant for every fixed $n$)
$\Upsilon(u)(n)=x\quad\text{ if }\quad\Upsilon(a)(n)=x\text{ for all $a\in[u]$}.$
Consider the tree $T$ of all triplets $(u,v,s)\in\bigcup_{m}(\prod_{j<m}{\mathcal{M}}_{j})^{2}\times{\mathcal{P}}(m)$ such that all $j\in s$ satisfy $u(j)=v(j)$, but
[…]
ordered by the end-extension. By the assumption on $\Upsilon$, in $T$ there is no branch $t_{i}=(u_{i},v_{i},s_{i})$ such that $\bigcup s_{i}$ is infinite. Fix a node $t=(u,v,s)\in T$ with the property that if $t^{\prime}=(u^{\prime},v^{\prime},s^{\prime})\in T$ extends $t$ then $s^{\prime}\setminus s$ is empty, and let $m$ be such that $u\in\prod_{j<m}{\mathcal{M}}_{j}$. We claim that for all $a$ and $b$ in $[u]$ and all $j\geq m$, if $a(j)=b(j)$ then $\Upsilon(a)(j)=\Upsilon(b)(j)$. Assume otherwise and fix $c\in[v]$ such that $c(j)=a(j)$. Then $\Upsilon(a)(j)=\Upsilon(c)(j)=\Upsilon(b)(j)$; contradiction.
For $j\geq m$ define $h_{j}\colon{\mathcal{M}}_{j}\to{\mathcal{N}}_{j}$ by $h_{j}(x)=\Upsilon(a)(j)$ for any $a\in[u]$ such that $a(j)=x$. These functions are as required.
This completes the proof in case when all ${\mathcal{M}}_{n}$ are finite. If this is not the case, let ${\mathcal{X}}$ be the family of all sets of the form $F=\prod_{n}F_{n}$, where $F_{n}$ is a finite subset of ${\mathcal{M}}_{n}$ for all $n$. By the first part of the proof, for each $F\in{\mathcal{X}}$ there are functions $h^{F}_{n}:F_{n}\to{\mathcal{N}}_{n}$, for $n\in{\mathbb{N}}$, such that $\forall^{\infty}n\ \Theta(a)(n)=h_{n}^{F}(a(n))$ for all $a\in F$.
On ${\mathcal{X}}$ consider the ordering $F\preceq^{*}G$ if $(\forall n)(F_{n}\subseteq G_{n})$. We may assume that the set $A=\{n\in{\mathbb{N}}\mid{\mathcal{M}}_{n}$ is infinite}}\}} is infinite, since otherwise we may assume that it is empty and the first part of the proof shows that in this case the conclusion holds. The set of those $F\in{\mathcal{X}}$ such that $F_{n}={\mathcal{M}}_{n}$ for all $n\in{\mathbb{N}}\setminus A$ is clearly cofinal in ${\mathcal{X}}$. The first part of the proof provides us with $h_{n}$, for $n\in{\mathbb{N}}\setminus A$, as required and we may therefore assume $A={\mathbb{N}}$.
Claim 5.11.
For every uncountable ${\mathcal{H}}\subseteq{\mathcal{X}}$ there exists $G\in{\mathcal{X}}$ such that the set $\{F\in{\mathcal{H}}\mid F\preceq^{*}G\}$ is uncountable.
Proof.
For each $n$ fix an enumeration of ${\mathcal{M}}_{n}$ in type $\omega$. Let ${\mathcal{X}}_{1}$ be the set of all $F\in{\mathcal{X}}$ such that $F_{n}$ is an initial segment of ${\mathcal{M}}_{n}$. Then ${\mathcal{X}}_{1}$ is cofinal in ${\mathcal{X}}$ and clearly $({\mathcal{X}}_{1},\preceq^{*})$ is isomorphic to $({\mathbb{N}}^{{\mathbb{N}}},\leq^{*})$, in particular $\operatorname{{\mathsf{OCA}_{T}}}$ implies that every subset of ${\mathcal{X}}_{1}$ of cardinality $\aleph_{2}$ is $\preceq^{*}$-bounded, because it implies $\mathfrak{b}=\aleph_{2}$. The claim therefore follows. ∎
Let $[{\mathcal{X}}]^{2}=K_{0}\cup K_{1}$ be defined by letting $\{F,G\}\in K_{0}$ if there are $n$ and $x\in F_{n}\cap G_{n}$ such that $h^{F}_{n}(x)\neq h^{G}_{n}(x)$. This is an open partition in the separable metric topology obtained by identifying $F\in{\mathcal{X}}$ with the pair $(F,(h^{F}_{n}:n\in{\mathbb{N}}))$.
Suppose that ${\mathcal{X}}$ has an uncountable $K_{0}$-homogeneous set, ${\mathcal{H}}$. By Claim 5.11 there is $G\in{\mathcal{X}}$ such that =${\mathcal{H}}’=\{F\in{\mathcal{H}}\mid F_{n}\subseteq G_{n}$ for all $n\}$ is uncountable. For every $F\in{\mathcal{H}}$ there is $m(F)$ such that for all $j\geq m(F)$ we have $h^{G}_{j}\restriction F_{j}=h^{F}_{j}$. Fix $m$ such that ′′=${\mathcal{H}}^{\prime\prime}=\{F\in{\mathcal{H}}^{\prime}\mid m(F)=m\}$ is uncountable. Since the set $\bar{G}=\prod_{j<m}G_{j}$ is finite, there are only countably many possibilities for the restriction of $h$ to this set. Therefore there are $F$ and $F^{\prime}$ in ${\mathcal{H}}^{\prime\prime}$ such that $h_{j}^{F}=h_{j}^{F^{\prime}}$ for all $j<m$. This implies $\{F,F^{\prime}\}\in K_{1}$; contradiction.
Therefore $\operatorname{{\mathsf{OCA}_{T}}}$ implies that ${\mathcal{X}}$ can be covered by countably many $K_{1}$-homogeneous sets. Consider the order on ${\mathcal{X}}$ defined by $F\leq^{*}G$ if $F_{j}\subseteq G_{j}$ for all but finitely many $j$. This order is $\sigma$-directed, and therefore one of the $K_{1}$-homogeneous sets, ${\mathcal{Y}}$, is $\leq^{*}$-cofinal in ${\mathcal{X}}$. Moreover there is $m\in{\mathbb{N}}$ such that for every $F\in{\mathcal{X}}$ there is $G\in{\mathcal{Y}}$ such that $F_{j}\subseteq G_{j}$ for all $j\geq m$. Then if $j\geq m$, $h_{j}=\bigcup_{F\in{\mathcal{Y}}}h^{F}_{j}$ is a function from ${\mathcal{M}}_{j}$ into ${\mathcal{N}}_{j}$. If $j<m$ fix an arbitrary $h_{j}$. Then for every $a\in\prod_{n}{\mathcal{M}}_{n}$ there are $F\in{\mathcal{Y}}$ and $b\in F$ such that $a(j)=b(j)$ for all $j\geq m$, and therefore (using the notation from Definition  2.1)
$\Phi[h_{n}](\pi(a))=\Phi[h_{n}^{F}](\pi(b))=\pi(\Upsilon(b))=\pi(\Upsilon(a)).\qed$
It is very likely that the version of Proposition 5.10 in which $\Theta$ is assumed to be Borel-measurable can be proven in $\operatorname{{\mathsf{ZFC}}}$, using the fact that $\operatorname{{\mathsf{OCA}_{T}}}$ for analytic sets is a theorem of $\operatorname{{\mathsf{ZFC}}}$ ([17]). A similar argument was used in [9, Theorem 1.6.1 and 1.6.2].

### v6

@@ 5.2. Uniformization
Assume $\operatorname{{\mathsf{OCA}_{T}}}$. Let ${\mathcal{M}}_{n}$ and ${\mathcal{N}}_{n}$, for $n\in{\mathbb{N}}$, be countable sets, and let
@@
be a coordinate-respecting function with $\alpha^{\Phi}=\operatorname{id}$ such that the ideal ${\mathcal{J}}_{\rm{prod}}$ intersects every perfect tree-like almost disjoint family nontrivially. Then $\Phi$ is of product form.
@@
Using the notation set up before Definition 2.6, if $A\in\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$, then $\Phi_{A}$ is the induced function between $\prod_{n}{\mathcal{M}}_{n}/\operatorname{Fin}\restriction A$ and $\prod_{n\in A}{\mathcal{N}}_{n}/\operatorname{Fin}\restriction A$.
@@
of functions with $g^{A}_{n}\colon{\mathcal{M}}_{n}\to{\mathcal{N}}_{n}$ such that the function $\Psi_{A}\colon\prod_{n\in A}{\mathcal{M}}_{n}\to\prod_{n\in A}{\mathcal{N}}_{n}$ defined by
@@
Consider the Polish space ${\mathcal{Y}}_{n}=({\mathcal{N}}_{n})^{{\mathcal{M}}_{n}}$. Then $\{(g^{A},A)\mid A\in{\mathcal{J}}_{\rm{prod}}\}$ is a family of partial selectors as defined and studied in §3.1. The following shows that the functions $g^{A}$, for $A\in{\mathcal{J}}_{\rm{prod}}$, form a coherent family in the sense of [10, §2].
@@
By assumption, ${\mathcal{J}}_{\rm{prod}}$ intersects every perfect tree-like almost disjoint family. Fix $B\in{\mathcal{A}}\cap{\mathcal{J}}_{\rm{prod}}$. For $x\in Z$ let $n(x)=\max(\operatorname{Diff}(g^{f(x)},g^{B}))$ (with $\max\emptyset=0$). Since $Z$ is uncountable, there is $n$ such that $Z^{\prime}=\{x\in Z\mid n(x)<n\}$ is uncountable. Choose distinct $y,z$ in $Z^{\prime}$ satisfying $\Delta(y,z)\geq n$ and fix $j$ in $(\operatorname{Diff}(g^{f(y)},g^{f(z)})\cap B)\setminus n$. Then $g^{B}_{j}=g^{f(y)}_{j}\neq g^{f(z)}_{j}=g^{B}_{j}$; contradiction. ∎
Since ${\mathcal{J}}_{\rm{prod}}$ is nonmeager (again by Talagrand and Jalali-Naini, [10, Theorem 3.10.1]), there is $\bar{n}$ such that ${\mathcal{G}}_{\bar{n}}$ as provided by Claim 5.8 is nonmeager. We will attempt to find a sequence $(A(i),B(i),k(i),x(i))$, for $i\in{\mathbb{N}}$, with the following properties for all $i$.
  1. (1)
    $A(i)$ and $B(i)$ belong to ${\mathcal{G}}_{\bar{n}}$.
  2. (2)
    $k(i)\in A(i)\cap B(i)$.
  3. (3)
    $x(i)\in{\mathcal{M}}_{k(i)}$.
  4. (4)
    $g^{A(i)}_{k(i)}(x(i))\neq g^{B(i)}_{k(i)}(x(i))$.
  5. (5)
    If $1\leq j<i$ then $g^{A(i)}_{k(j)}(x(j))=g^{A(j)}_{k(j)}(x(j))$ and $g^{B(i)}_{k(j)}(x(j))=g^{B(j)}_{k(j)}(x(j))$.
  6. (6)
    Each of the following two sets is nonmeager.
    ${\mathcal{A}}_{i}$ =$=\{C\in{\mathcal{G}}_{\bar{n}}\mid(\forall j<i)k(j)\in C\text{ and }g^{C}_{k(j)}(x(j))=g^{A(j)}_{k(j)}(x(j))\},$
    ${\mathcal{B}}_{i}$ =$=\{C\in{\mathcal{G}}_{\bar{n}}\mid(\forall j<i)k(j)\in C\text{ and }g^{C}_{k(j)}(x(j))=g^{B(j)}_{k(j)}(x(j))\}.$
These conditions imply that $\operatorname{Diff}(g^{A(i)},g^{B(i)})\supseteq\{k(j)\mid j\leq i\}$, and by the choice of ${\mathcal{G}}_{\bar{n}}$, the recursive construction of this sequence has to stop before $i=\bar{n}$. Note that ${\mathcal{A}}_{0}={\mathcal{B}}_{0}={\mathcal{G}}_{\bar{n}}$ and set $k(-1)=-1$.
Fix the least $i\leq\bar{n}$ such that $(A(i),B(i),k(i),x(i))$ cannot be chosen (such $i$ exists and $i<\bar{n}$). What this implies is that for every $k>k(i-1)$, every $x\in{\mathcal{M}}_{k}$, and all $y\neq z$ in ${\mathcal{N}}_{k}$, at least one of the sets $\{C\in{\mathcal{A}}_{i}\mid g^{C}_{k}(x)=y\}$ and $\{C\in{\mathcal{B}}_{i}\mid g^{C}_{k}(x)=z\}$ is meager. Since ${\mathcal{M}}_{k}$ and ${\mathcal{N}}_{k}$ are countable, this means that for every $k>k(i-1)$ there is a function $\tilde{g}_{k}\colon{\mathcal{M}}_{k}\to{\mathcal{N}}_{k}$ such that the set
${\mathcal{H}}_{k}=\{C\in{\mathcal{A}}_{i}\cup{\mathcal{B}}_{i}\mid g^{C}_{k}=\tilde{g}_{k}\}$
is relatively comeager in ${\mathcal{A}}_{i}\cup{\mathcal{B}}_{i}$. Therefore $\bigcap_{k>k(i-1)}{\mathcal{H}}_{k}$ is relatively comeager in ${\mathcal{A}}_{i}\cap{\mathcal{B}}_{i}$. Let $\hat{\mathcal{H}}$ denote the hereditary closure of this set.
For $A\in\hat{\mathcal{H}}$ pick $C\in{\mathcal{H}}$ such that $A\subseteq B$ and let $g^{A}_{k}=g^{C}_{k}$ for all $k\in A\setminus k(i-1)$. Note that $g^{A}_{k}=\tilde{g}_{k}$ for all $k\in A\setminus k(i-1)$. [10, Theorem 3.10.1(a)] implies that there exists $\bar{k}$ such that every finite subset of ${\mathbb{N}}\setminus\bar{k}$ is in $\hat{\mathcal{H}}$.
@@
For every $a\in\prod_{n}{\mathcal{M}}_{n}$ the set $S_{a}=\{k\mid\Phi_{*}(a)(k)\neq\tilde{g}_{k}(a(k))\}$ is finite.
@@
Assume otherwise. Since $\hat{\mathcal{G}}_{\bar{n}}$ is nonmeager and hereditary, there is an infinite $X\subseteq S_{a}$ in $\hat{\mathcal{H}}$ ([10, Theorem 3.10.1 (a)] gives a considerably stronger statement). Let $Y\in{\mathcal{H}}$ be such that $X\subseteq Y$. We have $g^{Y}_{k}=\tilde{g}_{k}$ for all $k>k(i-1)$, in particular for all $k\in X\setminus k(i-1)$. Therefore $g^{Y}_{k}(a(k))\neq\Phi_{*}(a)(k)$ for infinitely many $k$, contradicting our assumption that $g^{Y}$ witnesses $Y\in{\mathcal{J}}_{\rm{prod}}$. ∎
Claim 5.9 implies that $\tilde{g}_{k}$, for $k>k(i-1)$, provide a lifting of product form. This is the thesis. ∎
@@
Remark 5.10.
In a previous version of this manuscript we had a different, longer, proof of Theorem 5.1. We had shown that the hypotheses of Proposition 5.6 imply the existence of a Baire-measurable lifting and that every coordinate-respecting map between reduced products with Baire-measurable lifting is of twisted product form.
Remark 5.11.
In the case when $\sup_{n}|{\mathcal{M}}_{n}|$ is finite, a considerably simpler argument shows that every coordinate-respecting map $\Phi\colon\prod{\mathcal{M}}_{n}/\operatorname{Fin}\to\prod{\mathcal{N}}_{n}/\operatorname{Fin}$ such that $\alpha^{\Phi}=\operatorname{id}$ is automatically of product form. Indeed, suppose that $\sup_{n}|{\mathcal{M}}_{n}|=m$ is finite and fix enumerations ${\mathcal{M}}_{n}=\{x^{n}_{i}:i<|{\mathcal{M}}_{n}|\}$ for each $n$. For $i<m$, define $a_{i}\in\prod_{n}{\mathcal{M}}_{n}$ by
$a_{i}(n)=\begin{cases}x^{n}_{i}&\text{if }i<|{\mathcal{M}}_{n}|,\\ x^{n}_{0}&\text{otherwise}.\end{cases}$
It is easy to verify that the sequence $(g_{n})_{n}$ of maps $g_{n}\colon{\mathcal{M}}_{n}\to{\mathcal{N}}_{n}$ defined by $g_{n}(x^{n}_{i})=\Phi_{\ast}(a_{i})(n)$ satisfies $\Phi[(g_{n})]=\Phi$. In this case Propositions 5.2 and  5.6 are a fortiori true in $\operatorname{{\mathsf{ZFC}}}$, and Theorem 5.1 follows from (in fact, is equivalent to) the statement that every automorphism of the Boolean algebra ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ is trivial.
We suspect that this equivalence does not hold in general, when the structures ${\mathcal{M}}_{n}$ are allowed to be infinite.
The assumption that $\alpha^{\Phi}=\operatorname{id}$ cannot be dropped from the above argument. If ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ is identified with $\prod_{\operatorname{Fin}}\{0,1\}$ (where $\{0,1\}$ is the two-element Boolean algebra), then each one of its automorphisms is coordinate-respecting, yet $\operatorname{\mathrm{CH}}$ implies that some automorphisms are nontrivial, or equivalently, not of product form.